OpenStudy (anonymous):

Evaluate the definite integral from 1 to 2 of x/(x+1)^2

OpenStudy (anonymous):

let u = x +1, so du = dx x = u - 1 Write the whole thing in terms of u and you should be able to integrate it.

OpenStudy (anonymous):

It would be something like this \[\int\limits_{2}^{3} (u -1)/u^2\] Forget to tell you to change the limits.

OpenStudy (anonymous):

Thank you for the reply. I understand the u-substitution part. I don't understand the change of limits though. What do we do that for?

OpenStudy (anonymous):

u = x +1

OpenStudy (anonymous):

plug in the value for x, in the upper limit you had x = 1 so u = 1 + 1 = 2. Same for the lower limit.